# HumanVBench: Exploring Human-Centric Video Understanding Capabilities of MLLMs with Synthetic Benchmark Data

### v2

@@ 3.3 Distractor-Included QA Generation Pipeline
For tasks with exact answers (e.g., restricted categories, numbers, or letters), questions, correct answers, and distractors are constructed using tailored templates and annotations from Section 3.2. For open-ended questions, we developed a pipeline to generate questions, answers, and distractors, applied to six tasks in HumanVBench: Human Emotion Recognition, Emotion Temporal Analysis, Emotion Intensity Comparison, Human-to-Text, Behavior Temporal Analysis, and Behavior Causality Analysis. The Distractor-Included QA Synthesis Pipeline is illustrated in Figure 3.
3.3.1 Selecting Videos for Generating Questions
For a specific task, annotated videos are filtered based on task-specific criteria, such as video length and the number of people, to select suitable candidates for question generation. For instance, the Behavior Temporal Analysis task requires longer videos, while Human-to-Text necessitates at least two people. This process yields a subset of videos tailored for question generation.
Models Frames ER ETA AR EIC Avg T2H H2T HC ATD Avg BTA BCA AST TSA Avg AVSM ASD AVAD SCM Avg
Random 25 25.0 25.0 25.0 22.9 27.9 25.0 23.1 25.0 25.3 25.0 25.0 25.0 20.0 23.8 42.8 23.6 33.3 25.0 31.2
Chat-UniVi 1 f/s 29.2 19.9 10.9 16.3 18.8 29.1 37.2 17.8 19.8 26.0 22.0 19.4 14.0 6.7 15.5 42.8 20.6 26.3 18.6 27.1
CogVLM2-Video 1 f/s 25.0 33.1 36.0 11.1 25.8 42.5 63.8 31.1 40.0 44.4 42.4 47.8 34.0 14.1 34.6 59.2 38.7 30.7 17.9 36.6
VideoLLaMA3 1 f/s 39.4 34.5 48.0 43.1 41.3 90.3 78.1 44.9 71.0 71.1 72.4 53.4 56.4 69.0 62.8 64.6 63.1 34.4 17.9 45.0
VILA 6 f 29.1 34.4 28.7 14.7 26.7 50.4 54.2 40.4 20.2 41.3 53.0 54.5 33.0 51.4 48.0 47.4 23.9 34.3 18.6 31.1
Video-LLaVA 8 f 27.1 25.7 26.2 24.7 25.4 27.9 46.8 28.3 31.7 33.7 34.8 37.3 27.0 40.7 35.0 50.0 28.4 34.3 21.2 33.5
LLaVAOneVision 8 f 36.5 33.1 62.8 15.7 36.0 67.0 68.1 49.1 10.0 48.6 63.6 55.2 37.0 47.5 50.8 52.6 51.6 35.0 26.9 42.8
InternVL2 8 f 35.4 29.3 40.2 25.9 32.7 70.4 59.6 37.2 20.8 47.0 58.5 52.2 38.0 33.8 45.6 51.7 55.0 33.6 25.5 41.5
InternVL2.5 8 f 44.8 31.6 54.3 51.5 45.6 81.0 78.7 40.6 35.8 59.0 72.9 55.2 41.0 62.1 57.8 65.1 61.3 32.1 15.4 43.5
Qwen-VL2 8 f 41.7 40.4 42.7 32.8 39.1 79.3 69.2 43.4 20.8 53.2 61.9 55.2 32.0 51.4 50.1 50.7 56.1 31.4 23.7 40.5
Qwen-VL2.5 8 f 43.7 31.6 35.4 46.5 39.3 88.8 77.6 50.9 30.8 62.0 72.9 47.8 34.0 50.8 51.4 71.0 61.3 33.6 18.8 46.2
PLLaVA 16 f 25.0 24.3 36.6 23.2 26.5 33.5 54.3 29.9 23.3 35.3 30.5 32.8 34.0 14.7 28.0 50.7 37.4 29.9 21.8 35.0
ShareGPT4Video 16 f 36.5 30.9 40.9 10.1 29.6 33.0 40.4 24.5 43.3 35.3 45.8 40.3 39.0 16.9 35.5 44.1 31.0 34.3 25.0 33.6
Otter-V 16 f 15.6 25.7 26.2 21.2 21.9 30.2 22.3 20.8 22.5 24.0 22.9 17.9 20.0 23.1 21.0 38.2 26.5 33.6 27.6 31.5
VideoChat2-IT 16 f 25.0 30.9 39.6 25.3 30.2 20.1 47.9 11.2 26.7 26.5 42.4 47.8 34.0 24.9 37.3 43.4 27.7 31.4 23.0 31.4
LLaVA-Video 64 f 40.6 32.4 59.8 37.4 41.8 74.3 72.3 44.3 26.7 54.4 68.6 52.2 47.0 58.8 56.7 52.0 58.7 32.8 39.3 45.7
Video-LLaMA 8 f 25.0 25.9 11.5 19.7 18.9 28.4 33.2 23.9 20.0 26.4 39.5 32.5 23.2 15.4 27.7 40.1 26.6 33.1 26.2 31.5
VideoLLaMA2.1-AV 8 f 32.3 28.7 47.0 21.7 31.9 34.6 47.9 41.5 16.7 35.2 47.5 43.3 27.0 33.3 37.8 44.0 31.6 32.1 23.7 32.9
ImageBind-LLM 15 f 15.6 21.0 25.0 26.4 21.0 23.5 24.5 23.8 19.5 22.8 19.5 19.4 24.0 23.3 21.6 45.0 25.1 28.9 22.9 30.5
ChatBridge 4 f 27.1 15.6 30.2 14.4 21.8 31.4 41.5 23.7 8.7 26.3 31.4 23.9 41.0 16.9 28.3 43.7 25.3 30.4 24.4 30.9
OneLLM 15 f 26.0 29.4 34.1 21.1 26.1 29.0 38.3 20.6 27.8 28.9 32.2 38.8 23.0 22.4 29.1 43.4 26.4 29.5 23.2 30.6
GPT-4o 54.1 38.2 27.4 64.1 40.1 46.9 83.0 47.2 32.5 50.7 81.4 67.2 47.1 61.6 64.3 - - - - -
Gemini-1.5-Pro 57.3 54.4 53.0 67.2 55.9 87.1 77.7 52.8 71.7 72.3 78.0 65.7 54.0 75.1 68.2 90.1 76.8 66.4 84.6 79.5
Human 91.6 86.0 87.8 80.8 86.6 98.9 85.8 92.5 78.3 88.9 93.0 86.5 88.6 88.1 89.1 96.0 96.1 87.0 88.0 91.8
Table 1: The performance of SOTA video MLLMs on HumanVBench, grouped into 15 visual-only MLLMs, 5 audio-visual MLLMs and 2 proprietary ones. “Random” denotes random guessing, and “Human” indicates human-level performance. Task acronyms are defined in Figure 1. For each task, the best overall result is bolded, and the best open-source result is underlined. “-” means the model recognizes its lack of required capabilities for the task and thus refuses to answer. The results for all open-source models are averaged over five runs with different random seeds. Additional evaluation details are in the Appendix Section 9.
3.3.2 Question and Preliminary Answer Synthesis
This stage synthesizes questions and initial answers by processing the video with bounding box track information of the target person. For tasks focusing on facial expressions (Human Emotion Recognition and Emotion Temporal Analysis), we reconstruct the video using cropped face regions; for others, we add a red bounding box to create “marked videos”. Then, we design prompts for Video-MLLM to generate task-specific captions centered on the target individual. We find that video-MLLMs don’t always focus correctly on the person highlighted by the bounding box. To address this, we enhance attention by adding a visual cue of the target person’s appearance from Label-4 (Figure 2).
After acquiring task-specific captions, we analyze the caption attribute distributions (e.g., sentiment distribution for emotion recognition) and filter materials for question generation based on the desired distribution. Using these captions, we leverage SOTA LLMs to formulate questions and preliminary answers according to task definitions.
3.3.3 Answer Optimization and Distractors Generation
To obtain the most suitable answer, we iteratively optimize the answer using three Video-MLLMs while generating three distractors. Each Video-MLLM first generates a candidate answer from the video and question, then selects a better answer between this candidate and the current optimal answer (initialized as the preceding preliminary answer), updating the optimal answer. The discarded answers serve as raw material for distractor generation. This process repeats three times, producing one optimal answer and three discarded answers. An LLM then modifies the discarded answers with task-specific disturbances, ensuring distractors are distinct from the correct answer.
By using MLLM-discarded answers to synthesize options, the distractors are more challenging compared to those directly generated from the question and answer, thus more rigorously testing the model’s discriminative ability.
3.3.4 Manual Verification and Correction
Given the limitations of the annotation and question-answer generation models, errors in the multiple-choice questions may occur. To address this, we conduct manual validation as a final step. Specifically, we shuffle the options of each generated question and ask verifiers to select the best answer. If their choice matches the pipeline’s automatic selection, we confirm the question aligns with human cognition. If not, we re-examine and determine whether to revise the correct answer. If no suitable option is found, verifiers provide the correct answer, which becomes the definitive answer for the question. While this process involves some manual effort, it is significantly less labor-intensive and costly than generating questions and answers from scratch.
The Distractor-Included QA Synthesis Pipeline automates the tedious process of creating questions and options, shifting the human role from question creators to quality inspectors. Notably, we’ve decomposed the generation process into simpler sub-tasks, each manageable by open-source (M)LLMs. While these models are less accurate than their commercial counterparts, their inherent noise enhances the cognitive complexity of distractors. After human review and correction, this approach can still ensure high question quality while reducing labor costs. The pipeline is highly transferable, enabling its adaptation for designing a broader range of descriptive questions and choices.
@@ 3.4 Post-Processing
To address the prevalent issue of answer leakage [10] in multimodal evaluation datasets, we adopt the approach proposed in [10]. Specifically, we test all evaluated models without visual input and remove frequently correct QAs (approximately 6%) to ensure random accuracy. This strategy preserves the visual relevance of the questions while effectively mitigating the risk of answer leakage. Finally, HumanVBench includes 2116 problem instances. Instance counts for each task are detailed in Figure 1, and additional statistics are available in Appendix Section 7.
Figure 4: Two examples of 8-frame speaker videos sampled at equal intervals in the emotion recognition task, along with the responses from different MLLMs.
Models Chat-UniVi LLaVA-OneVision Intern VL2 Share-GPT4-Video Video Chat2-IT Qwen-VL2 LLaVA-Video Video-LLaMA Video-LLaMA-2.1 Chat Bridge GPT-4o Gemini-1.5-Pro Avg. Dec.
Emo. Acc. 29.2 36.5 35.4 36.5 25.0 41.7 40.6 25.0 32.3 27.1 54.1 57.3 36.7
Speaker Emo. Acc. 24.3$\downarrow$-4.9 34.3$\downarrow$-2.2 27.1$\downarrow$-8.3 32.8$\downarrow$-3.7 21.4$\downarrow$-3.6 37.1$\downarrow$-4.6 37.1$\downarrow$-3.5 22.9$\downarrow$-2.1 31.4$\downarrow$-0.9 21.4$\downarrow$-5.7 51.6$\downarrow$-2.5 54.0$\downarrow$-3.3 32.9$\downarrow$-3.8
Table 2: Accuracy of Video-MLLMs on Emotion Recognition (ER) for the full dataset and the subset where target individuals are in the speaking state. Current video-MLLMs struggle with emotion recognition for speaking individuals in HumanVBench.
@@ 4.1 Experimental Settings
We meticulously select 20 SOTA open-source video MLLMs. These included both visual-only models such as Chat-UniVi [25], CogVLM2-video [21], LLaVA-One-Vision [28], PLLaVA-7b [53], ShareGPT4Video [11], Otter-V [27], VILA [35], VideoChat [29], InternVL series [14, 15], Qwen-VL series [48, 4] and VideoLLaMA3 [57] and audio-visual models capable of analyzing both visual and audio inputs, such as the Video-LLaMA series [58, 16] and generalist MLLMs like ImageBind-LLM [19], ChatBridge [61], and OneLLM [20]. We also evaluated commercial models GPT-4o [41] and Gemini-1.5-Pro [47]. All QAs were framed as multiple-choice questions (N choose 1, with N varying across different test samples), reporting both accuracy and the performance of random guesses and graduate-level humans for reference. More implementation details can be found in the Appendix Section 9.
4.2 Main Results
@@ 4.2.1 Performance in Vision-Only Tasks
Open-Source Video-MLLMs: Generally, these models still show a noticeable gap compared to human-level performance, particularly in Emotion Perception tasks, highlighting the need for further development to bridge this gap. However, some models have demonstrated superior performance in certain tasks, rivaling or even surpassing commercial models, such as LLaVA-OneVision in AR and VideoLLaMA3 in T2H, ATD, and AST. Among open-source models, VideoLLaMA3 leads in video human understanding, and achieves a modest 58.4% mean accuracy across 12 vision-only tasks. Although it still lags behind human performance (88.2%), it has surpassed the commercial model GPT-4o (51.7%) and is close to Gemini-1.5-pro (65.5%), demonstrating the immense potential of open-source models.
Proprietary MLLMs: Overall, these models outperform their open-source counterparts in most tasks, achieving near-human proficiency in several tasks. Notably, Gemini delivers superior performance compared to other models, while GPT-4o struggles in tasks like AR, T2H, and ATD. As a result, GPT-4o’s average performance in Emotion Perception and Person Recognition lags behind several leading open-source models.
@@ 4.2.2 Performance in Speech-Visual Alignment Tasks
Lip-Reading Ability of Visual-Only MLLMs For AVSM and ASD tasks, many videos include dubbing or multi-speaker conversations with single-speaker audio, making lip movements alone insufficient for accurate responses. Therefore, when answering these two tasks, visual-only models essentially degrade to speech action recognition but can still get some questions correct. As shown in Table 1, these MLLMs show some lip-reading ability, with QwenVL2 and InternVL2 outperforming other models. However, for the SCM task, almost all models perform at a random level, indicating that current models lack precise lip-reading (lip translation) ability. Future models could focus on enhancing lip-reading capabilities.
Error Type Error Percentage
Misidentifying emotion as Surprise 41%
Overinterpretation of neutral emotions 29%
Emotion Polarity Mistakes 29%
Table 3: Error Analysis of Emotion Recognition in Gemini
Video-MLLMs ATD AST TSA
Chat-UniVi 17.5 ($\uparrow$2.3) 40 ($\uparrow$26) 14.1 ($\uparrow$7.4)
CogVLM2 40.8 ($\uparrow$0.8) 35 ($\downarrow$1) 10.7 ($\downarrow$3.4)
VILA 38.3 ($\uparrow$18.1) 33 (–) 48 ($\downarrow$3.4)
Video-LLaVA 36.7 ($\uparrow$5) 25 ($\downarrow$2) 45.8 ($\uparrow$5.1)
@@
InternVL2 31.7 ($\uparrow$10.9) 45 ($\uparrow$7) 52.5 ($\uparrow$18.7)
Video-LLaMA 24.2 ($\uparrow$4.2) 25 ($\uparrow$7) 9 (–)
Video-LLaMA2.1 24.2 ($\uparrow$7.5) 25 ($\downarrow$2) 29.4 ($\downarrow$3.9)
ChatBridge 17.5 ($\uparrow$8.8) 40 ($\downarrow$1) 14.1 ($\downarrow$2.8)
Average 26.8 ($\uparrow$5.9) 34.3 ($\uparrow$4.4) 30.1 ($\uparrow$2.0)
Table 4: Timestamp integration effect on Video-MLLMs in Appearance Time Detection (ATD), Action at Specific Time (AST) and Time of Specific Action (TSA).
Audio-Video MLLMs: Despite handling both audio and visual inputs, these open-source MLLMs perform near-randomly across Speech-Visual Alignment tasks, particularly in AVSM and ASD, where they lag behind visual-only MLLMs. On one hand, this may stem from deficiencies in the models’ ability to visually interpret lip movements (see Appendix Section 8 for extended experiments); on the other, contemporary open-source models largely neglect intricate speech-to-lip movement alignment. Few have explicit architectural encoding linking audio and video modalities to facilitate such synergy, compounded by a scarcity of comprehensive datasets proposing coherent visual-audio lexical mapping. Consequently, these models struggle acutely in correlating speech to lip movements, though dedicated ASR models effectively address Speech Content Matching — a capability absent in tested open-source video-MLLMs.
4.3 Discussion on Speaker Emotion Recognition
A detailed analysis reveals that models frequently misattribute speaker emotions as “surprise” or “shock.” For example, Gemini-pro-1.5, the top performer in Emotion Recognition, exhibits this tendency, with surprise misidentification accounting for 40% of its errors (Table 2). This issue primarily stems from frame sampling processes, where frames depicting “mouth-opening” motions are misclassified. Figure 4 illustrates this phenomenon using the commonly adopted eight-frame fixed sampling approach [34, 60, 28, 58], which introduces temporal noise and leads to erroneous emotional judgments. The speaker-centric evaluations presented in Table 2 further confirm a consistent decline in accuracy across all models, underscoring the inherent challenges of this task for video MLLMs.
4.4 Impact of Timestamp Inclusion on Time-Specific Tasks
Table 1 shows that many open-source video-MLLMs struggle with time-specific tasks—like Appearance Time Detection, Action at Specific Time, and Time of Specific Action—due to challenges in establishing “video scene-event time” correlations. Analysis shows that only a few models, including top performers like VideoLLaMA3 and LLaVA-Video, explicitly incorporate textual timestamps as part of their built-in prompt processing from sampled video frames. To explore the impact of timestamp integration, we added timestamps (e.g., “You have read N frames, corresponding to seconds in the video as: […]”) to models without native support. As shown in Table 4, this intervention generally improved temporal reasoning accuracy, though its effectiveness varied across models, likely due to limited exposure to timestamp-related data, reflecting the models’ inherent capabilities and potential.
@@ 4.4 Impact of Timestamp Inclusion on Time-Specific Tasks
Figure 5: Effectiveness in generating multiple-choice questions for six descriptive question types through the Distractor-Included QA Synthesis Pipeline.
4.5 Availability of Question Generation
The generation pipeline’s efficacy for multiple-choice questions across six descriptive categories is visualized in Figure 5. On average, 30% aligned directly with human responses, 20% included correct options conflicting with automated answers, and 50% required manual verification. While video-MLLMs and LLMs currently exhibit limitations in question generation, prospective advancements promise to enhance question accuracy, minimizing human intervention. Moreover, utilizing flawed video-MLLMs to craft misleading options augments evaluative complexity, thus refining model discriminative testing rigor.

### v3

@@ 3.3 Distractor-Included QA Generation Pipeline
For tasks with exact answers (e.g., restricted categories, numbers, or letters), questions, correct answers, and distractors are constructed using tailored templates and annotations from Section 3.2. For open-ended questions, we developed a pipeline to generate questions, answers, and distractors, applied to six tasks in HumanVBench: Emotion Recognition, Emotion Temporal Analysis, Emotion Intensity Comparison, Human-to-Text, Behavior Temporal Analysis, and Behavior Causality Analysis. The pipeline (Figure 3) consists of four stages:
Selecting Videos. Videos are filtered by task criteria. For example, Behavior Temporal Analysis requires longer clips, while Human-to-Text requires at least two people.
Question and Preliminary Answer Synthesis. Firstly, target persons are localized using bounding box tracks. For facial emotion-focused tasks, we reconstruct the video with cropped face regions; for others, the target person is highlighted with a red bounding box to create “marked videos”. These marked videos are fed into a Video-MLLM with prompt designed for the specific task, so that the model produces targeted descriptions such as emotional states or behavior summaries. Since Video-MLLMs may not always focus on the highlighted individual, we incorporate a visual cue from Label-4 (Figure 2) into the prompt. The generated descriptions are then analyzed to extract task-relevant attributes, and their distributions are adjusted to ensure balance (e.g., maintaining similar proportions of positive and negative emotions). Based on these captions, an LLM (GPT-4 in our workflow) is used to generate questions and preliminary answers.
Ensemble Answer Selection and Distractors Generation. To mitigate errors from single-model answers, we adopt an ensemble strategy. Multiple MLLMs (Gemini [47], VideoLLaMA3 [58], ShareGPT4Video [10]) produce candidate answers, which are ranked via preference voting. The top choice is taken as the correct answer, while the others are converted into distractors: if semantically distinct (valid answer component), they are retained; otherwise an LLM (GPT-4 in our workflow) introduces task-specific perturbations. Retaining these erroneous answers allows the distractors to reflect typical model mistakes, thereby ensuring plausibility and distractive power rather than being arbitrarily fabricated, which increases the overall difficulty of the questions.
Manual Verification and Correction Given the annotation models’ limitations, systematic biases and errors may occur in the generated multiple-choice questions. To address this concern and ensure the benchmark’s reliability, we incorporate a human verification stage: all candidate options are shuffled, and annotators are instructed to select the correct one; if none of the options is fully accurate, they directly write the correct answer, which is then incorporated as the ground truth.
The Distractor-Included QA Synthesis Pipeline handles large-scale question and distractor generation, while humans serve as final arbiters of correctness. This substantially reduces manual workload—annotators only needed to rewrite the correct answer in about 25% of cases, while in the remaining 75% they confirmed one of the generated options (see Section 4.4). To further ensure reliability, we conducted an inter-annotator agreement (IAA) study on 240 randomly sampled questions from the final verified questions, where two independent annotators achieved a Cohen’s Kappa of 0.8833, demonstrating the reliability of our benchmark. Additionally, we also provide distractor quality evaluation in Appendix E.
@@ 3.4 Post-Processing
To address the prevalent issue of answer leakage in multimodal evaluation datasets, we adopt the approach proposed in [9]. Specifically, we test models without visual input and remove frequently correct QAs (around 6%), ensuring random accuracy. This preserves visual relevance while mitigating leakage risks. In total, HumanVBench contains 2475 problem instances; per-task counts are shown in Figure 1, with additional statistics in Appendix C.
Table 1: Summary of performance on HumanVBench. We report average accuracy (%) across four main categories. Full results for all 30 models and 16 tasks are in Appendix D. The best overall results are shown in bold, while the best open-source model results are underlined “-” means the model recognizes its lack of required capabilities for the task and thus refuses to answer.
Model Mod. Emotion Person Recog. Behavior Avg. Speech-Visual Alignment Overall Avg.
4 tasks 4 tasks 4 tasks 12 tasks AVSM ASD AVAD SCM 4 tasks 16 tasks
Random Guess 24.4 25.2 22.9 24.2 42.8 23.6 33.3 25.0 31.2 25.9
Qwen-VL3 (7B) V 43.2 67.6 54.3 55.0 71.1 65.8 34.3 22.2 48.3 53.4
InternVL3.5 (8B) V 47.5 57.4 58.8 54.6 63.8 63.2 31.4 23.1 45.4 52.3
VideoLLaMA3 (7B) V 39.7 68.5 55.8 54.7 64.6 63.1 34.3 17.9 45.0 52.3
VideoLLaMA (7B) V+A 21.1 23.8 25.7 23.5 40.1 26.6 33.1 26.2 31.5 25.5
VideoLLaMA2.1 (7B) V+A 36.5 36.4 43.6 38.8 44.0 31.6 32.1 23.7 32.9 37.4
Qwen2.5-Omni (7B) V+A 35.5 44.5 38.3 39.4 71.1 48.4 27.0 71.8 54.6 43.2
GPT-4o V 33.6 50.9 62.1 48.9 - - - - - -
GPT-5 V 46.8 69.5 67.3 61.2 - - - - - -
Gemini-1.5-Pro V+A 50.9 71.4 60.7 61.0 90.1 76.8 66.4 84.6 79.5 65.6
Gemini-2.5-flash V+A 53.0 75.7 64.9 64.5 98.7 80.6 57.7 99.1 84.0 64.5
Gemini-2.5-Pro V+A 52.9 83.5 70.7 69.0 96.7 78.7 72.3 98.3 86.5 73.4
Human (Graduate) 84.6 88.5 87.0 86.7 96.0 96.1 87.0 98.3 94.4 88.6
@@ 4.1 Experimental Settings
We evaluate 30 of the most popular video MLLMs, including visual-only MLLMs from the Qwen-VL series, InternVL series, and VideoLLaMA3, as well as audio-visual MLLMs like Qwen2.5-Omni [52] and Video-LLaMA series [59, 15]. We also evaluated commercial models from the GPT-4o, GPT-5, and Gemini series. All tasks were cast as multiple-choice questions (N-choose-1, with variable N), with human and random guesses as baselines. For more implementation details and the full results of all 16 tasks on 30 MLLMs , please refer to Appendix I and D.
4.2 Comprehensive Evaluation of MLLMs
@@ 4.2.1 Performance in Vision-Only Tasks
Open-Source Video-MLLMs: Overall, current models still exhibit a clear gap from human-level performance, especially in Emotion Perception tasks, underscoring their limited ability to capture emotional nuances. Nevertheless, among open-source models, Qwen3-VL leads in video human understanding, and achieves a modest 55.0% mean accuracy across 12 vision-only tasks. Although it still lags behind human performance, it has surpassed GPT-4o and approaches other commercial counterparts, highlighting the immense potential of open-source models.
Proprietary MLLMs: Overall, with the exception of GPT-4o, proprietary models substantially outperform open-source ones across tasks, though their capability in emotion perception remains relatively weak, highlighting the need for better nuanced emotion understanding; notably, Gemini-2.5-Pro stands far ahead, achieving near-human proficiency in person recognition, whereas GPT-4o performs unexpectedly poorly; its subpar performance in the Emotion Understanding category and in specific Person Recognition tasks (T2H and ATD) results in its average performance being lower than that of several leading open-source models.
@@ 4.2.2 Performance in Speech-Visual Alignment Tasks
Lip-Reading Ability of Visual-Only MLLMs For AVSM and ASD tasks, many videos include dubbing or multi-speaker conversations with single-speaker audio, making lip movements alone insufficient for accurate responses. Therefore, when answering these two tasks, visual-only models essentially degrade to speech action recognition but can still get some questions correct, with QwenVL3 outperforming other models. However, for the AVAD and SCM task, almost all models perform at a random level, indicating that current models lack precise lip-reading (lip translation) ability. Future models could focus on enhancing lip-reading capabilities.
Audio-Video MLLMs: Although most open-source MLLMs can process both audio and visual inputs, they—except for Qwen2.5-Omni—perform near-randomly on Speech-Visual Alignment tasks, particularly AVSM and ASD, where they even lag behind visual-only models. This deficiency may stem from two key factors: a limited ability to visually interpret intricate lip movements (as further detailed in Appendix Section F) and a general lack of a suitable temporal processing mechanism to facilitate fine-grained speech-to-lip alignment, further exacerbated by scarce datasets providing audio-visual lexical mappings. Consequently, while dedicated ASR models can effectively handle SCM, most open-source video-MLLMs struggle acutely with correlating speech to lip movements. In contrast, the Gemini series shows exceptional cross-modal alignment, performing robustly even on AVAD, a task no other models can handle.
Figure 4: Two examples of 8-frame speaker videos sampled evenly for the emotion recognition task, with responses from various MLLMs.
4.3 Dissecting Fine-grained Insights
4.3.1 Discussion on Speaker Emotion Recognition
(a) Emotion Recognition Accuracy for the full dataset and the subset where target individuals are in the speaking state.
Models ER Accuracy
Full Speaking Subset
CogVLM2 38.4 34.7($\downarrow$3.7)
VideoLLaMA3 34.6 30.3($\downarrow$4.3)
LLaVAOneVision 36.9 34.7($\downarrow$2.2)
InternVL3.5 43.7 40.9($\downarrow$2.8)
InternVideo 35.7 31.0($\downarrow$4.7)
Qwen-VL2.5 43.0 41.0($\downarrow$2.0)
ShareGPT4Video 35.0 31.6($\downarrow$3.4)
ChatBridge 28.9 24.1($\downarrow$4.8)
Qwen2.5-Omni 42.2 40.3($\downarrow$1.9)
Average 36.9 33.5($\downarrow$3.3)
(b) Timestamp integration effect on Video-MLLMs in Appearance Time Detection , Action at Specific Time and Time of Specific Action.
Models Time-Specific Tasks
ATD AST TSA
Chat-UniVi 22.1 ($\uparrow$2) 40 ($\uparrow$26) 14.1 ($\uparrow$7)
CogVLM2 40.8 ($\uparrow$1) 35 ($\downarrow$1) 10.7 ($\downarrow$3)
VILA 38.3 ($\uparrow$18) 33 (–) 48 ($\downarrow$3)
Video-LLaVA 36.7 ($\uparrow$5) 25 ($\downarrow$2) 45.8 ($\uparrow$5)
@@ 4.3.1 Discussion on Speaker Emotion Recognition
InternVL2 31.7 ($\uparrow$11) 45 ($\uparrow$7) 52.5 ($\uparrow$19)
Video-LLaMA 24.2 ($\uparrow$4) 25 ($\uparrow$7) 9 (–)
Video-LLaMA2.1 24.2 ($\uparrow$7) 25 ($\downarrow$2) 29.4 ($\downarrow$4)
ChatBridge 17.5 ($\uparrow$9) 40 ($\downarrow$1) 14.1 ($\downarrow$3)
Average 26.8 ($\uparrow$6) 34.3 ($\uparrow$4) 30.1 ($\uparrow$2)
@@
(c) Effectiveness of the Distractor-Included QA Synthesis Pipeline in generating six types of descriptive multiple-choice questions. The Efficiency Gain is calculated as 1 / (1 - No Edit Rate).
Question Type No Edit Efficiency Gain
Behavior Temporal Analysis 68% 3.1
Emotion Intensity Comparison 63% 2.7
Emotion Recognition 81% 5.3
Behavior Causality Analysis 74% 3.8
Emotion Temporal Analysis 83% 5.9
Human-to-Text 65% 2.9
Average 72.3% 3.6
A detailed analysis reveals that models frequently misattribute speaker emotions as “surprise” or “shock”. This issue primarily stems from frame sampling processes, where frames depicting “mouth-opening” motions are misclassified. Figure 4 illustrates this phenomenon using the commonly adopted eight-frame fixed sampling approach [36, 62], which introduces temporal noise and leads to erroneous emotional judgments. The speaker-centric evaluations presented in Table 2(a) further confirm a consistent decline in accuracy across all models, underscoring the inherent challenges of this task for video MLLMs.
4.3.2 Timestamps Impact on Time-Specific Tasks
Our findings show that many open-source video-MLLMs struggle with time-specific tasks due to their inability to correlate video events with a timeline (Appendix Table D). While top performers like VideoLLaMA3 and LLaVA-Video natively incorporate textual timestamps, we explored this capability by adding timestamps to models without native support. As shown in Table 2(b), this intervention generally improves temporal inference accuracy, though its effectiveness varies due to models’ limited exposure to timestamp data. Despite these gains, a significant gap remains compared to MLLMs that natively incorporate time-event alignment, underscoring the necessity of integrating temporal reasoning mechanisms from the outset of model design.
4.4 Quantifying the Reduction of Manual Labor in Question Generation
We report the distribution of manual work required for the 6 descriptive multiple-choice questions generated by the pipeline in Table 2(c). On average, 72% of the generated questions contain the correct answer, requiring no revision, with humans only needing to select the correct answer. The remaining 28% lack a correct answer and require a simple manual input, typically one sentence, which is much faster than creating questions from scratch. Notably, the overall proportion of questions containing the correct answer exceeds the best model’s accuracy in the final evaluation (e.g., the “no-edit” rate for ER is 81%, while the best model accuracy is 49.4%), as options are drawn from multiple models, increasing correct answer coverage and reflecting the benefit of ensemble-based design. Besides, incorrect model responses are repurposed as plausible distractors, enhancing question difficulty. For other rule-based non-descriptive tasks, over 70% of the questions can be directly adopted. Overall, while human oversight is still needed, our partial automation significantly boosts efficiency, allowing a person using the pipeline to complete 3 times more questions within the same timeframe as manual creation from scratch, substantially reducing manual workload.
Figure 5: Demonstration of our pipeline’s generalizability. (a) Object attribute recognition for pets. (b) Object temporal tracking for cars. (c) General causal reasoning QA generated for an object-centric event, with the MLLMs used in this case being Gemini-1.5-pro, VideoLLaMA3, LLaVA-Video, and gpt-4o. This highlights the adaptability of our framework.
4.5 Generalizability of the Synthesis Pipelines Beyond Human-Centric Tasks
A key contribution of our work is the general-purpose nature of our automated pipelines, which extends far beyond the human-centric domain. Its modular architecture allows for easy adaptation to new domains by simply swapping initial detectors. We demonstrate this by generating diagnostic questions for non-human entities, such as fine-grained attribute recognition for pets and temporal tracking for vehicles, with examples shown in Figure 5.
For object attribute recognition (e.g., a specific dog; Figure 5 (a)), we adapt our pipeline by: (i) replacing the human detector with a dog detector (YOLOv11) to extract instance tracks (Label-1); (ii) using an MLLM to annotate attributes (e.g., fur pattern, accessories) from cropped tracks; (iii) selecting a uniquely described dog as the ground truth; and (iv) rendering bounding boxes from Label-1 as candidates. Similarly, for object temporal tracking (e.g., a car’s appearance interval, Figure 5 (b)), we: (i) replace the human detector with a vehicle detector for tracking; (ii) annotate attributes (e.g., color, type) for each track; (iii) select a unique car as ground truth; and (iv) use track frames to determine the appearance interval as ground truth. These procedures parallel our Text-to-Human and Appearance Time Detection tasks, proving the framework’s broad applicability to object-centric attribute recognition and temporal tracking.
Furthermore, our Distractor-Included QA Generation Pipeline is inherently transferable, enabling the synthesis of diverse descriptive questions and answer choices beyond human-centric scenarios, as exemplified by the general causal reasoning question in Figure 5 (c). This transferability underscores the broader value of our methodology as a scalable solution for accelerating the creation of high-quality MLLM evaluations across diverse domains.